# A Computational Theory for Efficient Mini Agent Evaluation with Causal Guarantees

### v4

@@ 1 Introduction
An improvement is to build the centralized A/B platform, which has been proven successes in application updates and recommendation in Byte Dance (Byte-Dance ). The implementation of centralized A/B testing platforms can significantly reduce the costs associated with conducting randomized trials for researchers. However, this cost reduction does not inherently enhance the utility derived from each individual trial and it still can not handle the infinite evaluation subjects problem. Notably, the lower cost and increased accessibility of such platforms can lead to a substantial rise in the number of randomized trials conducted. This proliferation of experiments raises ethical concerns regarding participant exposure to potential harm during the randomization process. For instance, as highlighted by Zhou et al. (2024b), the automation and scalability of A/B testing frameworks, while beneficial for rapid experimentation, necessitate careful consideration of ethical implications to safeguard participant well-being. Polonioli et al. (2023) emphasizes the need for ethical and responsible experimentation to protect users and society.
In addition to other agent evaluation works in appendix B, evaluatology (Zhan et al. (2024)) emerges as an independent filed in recent years. However, there is still a lack of a theory to help us find an effective and efficient evaluation system that satisfies the concerns and interests of the stakeholders. Here, we introduce a computational theory to evaluate the effect of given agents. The main benefit to evaluate by computation is its efficiency and low cost. If we learned the connection between the experiment result and pre-experiment agents, the evaluation cost will be reduced dramatically comparing with experiments.
@@ 2 Notation
Table 1 shows the notation we used in this paper. In order to simplify the evaluation problem, evaluation metric and difference of evaluation metric are assumed as measurable and pairwise comparable. The computational evaluation problem is modeled as computing the evaluation metrics for evaluation subject with certain evaluation conditions. Unless otherwise specified, all evaluation subjects in this paper are agent-based systems.
@@ 3.2 Hidden common cause
  • High-dimensional evaluation subject. Potential outcomes (Splawa-Neyman et al. (1990), Holland (1986), Imbens and Rubin (2015)) often assumes each evaluation subject must be sampled once which often do not established for high-dimensional evaluation subjects. For instance, parameters of AI model can be hundreds or thousands and more complex which structure is different from continuous dose-level analysis with very low dimension.
@@
Comparing with Saito and Yasui (2020) and Alaa and Van Der Schaar (2019), our unbiased assumption has potential to be rejected from real data testing as shown in the experiment part while they did not provide approaches to reject their unbiased causal effect assumption (Saito and Yasui (2020)) or min-max optimal effect estimator assumption (Alaa and Van Der Schaar (2019)) from real data without counterfactual. Also, IID samples assumption in those works (Saito and Yasui (2020), Johansson et al. (2022), Alaa and Van Der Schaar (2019), Shalit et al. (2017), Li and Pearl (2022)) is too strong for causal evaluation, especially IID conditions and IID outcomes.
@@ 6.2 Conditional evaluation model
Our best conditional evaluation models (HetEM-CatBoost) for heterogeneous subjects reduced 89.4% evaluation errors comparing with the baseline method. Even the linear conditional evaluation model can reduce the 34.0% RMSE of estimated RoI. The average RMSE of estimated RoI in a future time slot is about 0.527, which may be reduced further when adding more factors of stocks. From the result of different evaluation methods, the computational evaluation approach has higher potential to solve the backtest over-fitting problem (Bailey et al. (2013)) in stock market.
@@ 7 Limitation and conclusion
Limitation. First, our theoretical guarantees rely on three assumptions, which may not hold in all real-world scenes. Second, while our method has been validated across multiple domains, its generalization in extremely high-dimensional settings (such as LLM) require further investigation. Third, lower error and tighter bound are required to apply evaluation models into real applications with small sample. Four, scalability of evaluation models across scenes is another problem in future study. Five, the post-experiment metric (such as mortality, RoI) of randomized deployed mini-agents with high utility should be collected to increase the value of computational evaluation.
Conclusion. In this work, we propose a novel computational framework for the evaluation of mini-agents that rigorously derives upper bounds on generalized evaluation error and effect error, and employs a meta-learning strategy to address heterogeneity in agent space. Our extensive experimental results, conducted over 12 diverse scenes—including individual medicine, scientific simulation, social experiments, business activity, and quantum trade—demonstrate significant error reductions (ranging from 24.1% to 99.0%) and substantial acceleration (up to $10^{7}$) compared to traditional evaluation methods. These findings highlight the potential of our technical route to drastically reduce evaluation costs for rapid model iteration. Future work will focus on benefit-related metrics and enhancing the validation of key assumptions for applicability. The upper bound table and paradigm of computational evaluation were given in the appendix I and J.
@@ Appendix A Investigation of individual medical AI in randomized controlled trial
These results highlight the challenge of translating AI systems into demonstrably effective clinical interventions, emphasizing the need for rigorous evaluation, robust study design, and careful integration into complex healthcare workflows.
@@ Appendix C Proofs
. Because evaluation error is distributed on $[0,1]$, and Hoeffding inequality Hoeffding [1963],
@@ Appendix G Original evaluation errors
We listed the original average root mean square evaluation errors for 4 different evaluation metrics: ROC-AUC, ACC, RMSE, $R^{2}$ in the 11 scenes for evaluation model learning. The reason we do not use the policy risk, Area Under the Qini Curve (Qini Coefficient, Radcliffe [2007]) and Area Under the Uplift Curve (AUCC, Radcliffe [2007]) for discrete target is that not all models return a probability for classification. For example, the output continuous value of logistic regression can not be used to represent the probability that it belong to a class.
@@ Appendix I Upper bounds tables
We list the upper bound of generalized evaluation error and generalized causal effect evaluation error in the table 14 and table 15 for researchers’ quick query before data collection and upper bound estimation.
Samples/Confidence 0.5 0.8 0.95 0.99 0.999
10 $E+0.186$ $E+0.284$ $E+0.387$ $E+0.480$ $E+0.588$
20 $E+0.132$ $E+0.201$ $E+0.274$ $E+0.339$ $E+0.416$
30 $E+0.107$ $E+0.164$ $E+0.223$ $E+0.277$ $E+0.339$
100 $E+0.059$ $E+0.090$ $E+0.122$ $E+0.152$ $E+0.186$
1000 $E+0.0186$ $E+0.0284$ $E+0.0387$ $E+0.0480$ $E+0.0588$
10000 $E+0.00589$ $E+0.00897$ $E+0.0122$ $E+0.0151$ $E+0.0186$
100000 $E+0.00186$ $E+0.00284$ $E+0.00387$ $E+0.00480$ $E+0.00588$
1000000 $E+0.000589$ $E+0.000897$ $E+0.00122$ $E+0.00152$ $E+0.00186$
10000000 $E+0.000186$ $E+0.000284$ $E+0.000387$ $E+0.000480$ $E+0.000588$
100000000 $E+0.0000589$ $E+0.0000897$ $E+0.000122$ $E+0.000152$ $E+0.000186$
1000000000 $E+0.0000186$ $E+0.0000284$ $E+0.0000387$ $E+0.0000480$ $E+0.0000588$

### v5

@@ 1 Introduction
An improvement is to build the centralized A/B test platform, which has been proven successes in application updates and recommendation in Byte Dance (Byte-Dance ). The implementation of centralized A/B testing platforms can significantly reduce the costs associated with conducting randomized trials for researchers. However, this cost reduction does not inherently enhance the utility derived from each individual trial and it still can not handle the infinite evaluation subjects problem. Notably, the lower cost and increased accessibility of such platforms can lead to a substantial rise in the number of randomized trials conducted. This proliferation of experiments raises ethical concerns regarding participant exposure to potential harm during the randomization process. For instance, as highlighted by Zhou et al. (2024b), the automation and scalability of A/B testing frameworks, while beneficial for rapid experimentation, necessitate careful consideration of ethical implications to safeguard participant well-being. Polonioli et al. (2023) emphasizes the need for ethical and responsible experimentation to protect users and society.
In addition to other agent evaluation works in appendix B, evaluatology (Zhan et al. (2024)) emerges as an independent filed in recent years. However, there is still a lack of a theory to help us find an effective and efficient evaluation system that satisfies the concerns and interests of the stakeholders. Here, we introduce a computational theory to evaluate the effect of given agents. The main benefit to evaluate by computation is its efficiency and low cost. If we learned the connection between the experiment result and pre-experiment agents by a computational model, the evaluation cost will be reduced dramatically comparing with performing experiments.
@@ 2 Notation
Table 1 shows the notation we used in this paper. In order to simplify the evaluation problem, evaluation metric and difference of evaluation metric are assumed as measurable and pairwise comparable. The computational evaluation problem is modeled as computing the evaluation metrics for evaluation subject with certain evaluation conditions. Unless otherwise specified, all evaluation subjects in this paper are mini agents.
@@ 3.2 Hidden common cause
  • High-dimensional evaluation subject. Potential outcomes (Splawa-Neyman et al. (1990), Holland (1986), Imbens and Rubin (2015)) often assumes each evaluation subject must be sampled once which often do not established for high-dimensional evaluation subjects. For instance, parameters of AI model can be hundreds or thousands and more complex which structure is different from continuous dose-level analysis (Schwab et al. (2020), Nie et al. (2021), Zhu et al. (2024), Nagalapatti et al. (2024)) with low dimension.
@@
Comparing with Saito and Yasui (2020) and Alaa and Van Der Schaar (2019), our unbiased assumption has potential to be rejected from real data testing as shown in the experiment part while they did not provide approaches to reject their unbiased causal effect assumption (Saito and Yasui (2020)) or min-max optimal effect estimator assumption (Alaa and Van Der Schaar (2019)) from real data without counterfactual. Also, IID samples assumption in those works (Saito and Yasui (2020), Johansson et al. (2022), Alaa and Van Der Schaar (2019), Shalit et al. (2017), Li and Pearl (2022), Wang et al. (2022)) is too strong for causal evaluation, especially IID conditions and IID outcomes.
@@ 6.2 Conditional evaluation model
Our best conditional evaluation models (HetEM-CatBoost) for heterogeneous subjects reduced 89.4% evaluation errors comparing with the baseline method. Even the linear conditional evaluation model can reduce the 34.0% RMSE of estimated RoI. The average RMSE of estimated RoI in a future time slot is about 0.527, which may be reduced further when adding more factors of stocks, and designing novel neural network architectures. From the result of different evaluation methods, the computational evaluation approach has higher potential to solve the backtest over-fitting problem (Bailey et al. (2013)) in stock market.
@@ 7 Limitation and conclusion
Limitation. First, our theoretical guarantees rely on three assumptions, which may not hold for all real-world scenes and agents. Second, while our method has been validated across multiple domains, its generalization in extremely high-dimensional settings (such as LLM) require further investigation. Third, lower error and tighter bound are required to apply evaluation models into real applications with small sample. Four, scalability of evaluation models across scenes is another problem in future study. Five, the post-experiment metric (such as mortality) of randomized deployed mini-agents with high utility should be collected to increase the value of computational evaluation.
Conclusion. In this work, we propose a novel computational framework for the evaluation of mini-agents that rigorously derives upper bounds on generalized evaluation error and effect error, and employs a meta-learning strategy to address heterogeneity in agent space. Our extensive experimental results, conducted over 12 diverse scenes—including individual medicine, scientific simulation, social experiments, business activity, and quantum trade—demonstrate significant error reductions (ranging from 24.1% to 99.0%) and substantial acceleration (up to $10^{7}$) compared to traditional evaluation methods. These findings highlight the potential of our technical route to drastically reduce evaluation costs for rapid model iteration. Future work will focus on more benefit-related metrics, multi-source inputs, novel neural architectures for evaluation tasks and enhancing the validation of key assumption. The upper bound table and paradigm of computational evaluation were given in the appendix I and J.
@@ Appendix A Investigation of individual medical AI in randomized controlled trial
These results highlight the challenge of translating AI agent systems into demonstrably effective clinical interventions, emphasizing the need for rigorous evaluation, robust study design, and careful integration into complex healthcare workflows.
@@ Appendix C Proofs
. Because evaluation error is distributed on $[0,1]$, and Hoeffding inequality (Hoeffding [1963]),
@@ Appendix G Original evaluation errors
We listed the original average root mean square evaluation errors for 4 different evaluation metrics: ROC-AUC, ACC, RMSE, $R^{2}$ in the 11 scenes for evaluation model learning. The reason we do not use the policy risk, Area Under the Qini Curve (Qini Coefficient, Radcliffe [2007]) and Area Under the Uplift Curve (AUCC, Radcliffe [2007]) for discrete target is that not all models return a probability for classification. For example, the output continuous value of logistic regression can not be used to represent the probability that it belongs to a class.
@@ Appendix I Upper bounds tables
We list the upper bound of generalized evaluation error and generalized causal effect evaluation error in the table 14 and table 15 for researchers’ quick query before data collection and upper bound estimation after learning.
Samples/Confidence 0.5 0.95 0.99 0.999
10 $E+0.186$ $E+0.387$ $E+0.480$ $E+0.588$
20 $E+0.132$ $E+0.274$ $E+0.339$ $E+0.416$
30 $E+0.107$ $E+0.223$ $E+0.277$ $E+0.339$
100 $E+0.059$ $E+0.122$ $E+0.152$ $E+0.186$
1000 $E+0.0186$ $E+0.0387$ $E+0.0480$ $E+0.0588$
10000 $E+0.00589$ $E+0.0122$ $E+0.0151$ $E+0.0186$
100000 $E+0.00186$ $E+0.00387$ $E+0.00480$ $E+0.00588$
1000000 $E+0.000589$ $E+0.00122$ $E+0.00152$ $E+0.00186$
10000000 $E+0.000186$ $E+0.000387$ $E+0.000480$ $E+0.000588$
100000000 $E+0.0000589$ $E+0.000122$ $E+0.000152$ $E+0.000186$
1000000000 $E+0.0000186$ $E+0.0000387$ $E+0.0000480$ $E+0.0000588$